OpenStudy (idealist10):

Find the general solution of y"-3y'+2y=1/(1+e^(-x)), given that y1=e^2x satisfies the complementary equation.

OpenStudy (idealist10):

y=ue^(2x) y'=u'e^(2x)+2ue^(2x) y"=u"e^(2x)+4u'e^(2x)+4ue^(2x) Subbing: y"-3y'+2y=u"e^(2x)+u'e^(2x)=1/(1+e^(-x)) \[u''+u'=\frac{ e ^{-2x} }{ 1+e ^{-x} }\]

OpenStudy (idealist10):

z=u' subbing: \[z'+z=\frac{ e ^{-2x} }{ 1+e ^{-x} }\]

OpenStudy (idealist10):

the integrating factor is e^x \[e ^{x}(z'+z)=\frac{ e ^{-x} }{ 1+e ^{-x} }\]

OpenStudy (idealist10):

\[\frac{ d }{ dx }(e ^{x}z)=\frac{ e ^{-x} }{ 1+e ^{-x} }\] \[e ^{x}z=\int\limits_{}^{}\frac{ e ^{-x} }{ 1+e ^{-x} }dx\] \[e ^{x}z=-\ln \left| 1+e ^{-x} \right|+C\] \[z=u'=-e ^{-x}\ln \left| 1+e ^{-x} \right|+Ce ^{-x}\] \[u=(1+e ^{-x})\ln \left| 1+e ^{-x} \right|-(1+e ^{-x})+C _{2}e ^{-x}+C _{1}\] Since y=ue^(2x), \[y=(e ^{2x}+e ^{x})[\ln \left| 1+e ^{-x} \right|-1]+C _{1}e ^{2x}+C _{2}e ^{x}\]

OpenStudy (idealist10):

But the answer in the book says \[y=(e ^{2x}+e ^{x})\ln(1+e ^{-x})+C _{1}e ^{2x}+C _{2}e ^{x}\]

OpenStudy (idealist10):

@Kainui @freckles @pooja195

OpenStudy (freckles):

your answer is the same \[-e^{2x}-e^{x}+C_1 e^{2x}+C_2 e^x \\ =(C_1-1)e^{2x}+(C_2-1)e^{x} \\ =K_1 e^{2x}+K_2 e^{x}\]

OpenStudy (idealist10):

So my answer and the book's answer are both correct?

OpenStudy (freckles):

yep

OpenStudy (idealist10):

Okay, then. Thank you for verifying!